An empty maglev train car of mass M1= 24.2 metric tons and moving at a speed v1= 26.4 mph is hit from behind by another, heavier maglev car of mass M2= 44 metric tons and moving at a speed v2= 38.5 mph. Calculate the new speed, V (in m/s), of the system made of the two cars. Assume the collision to be perfectly elastic and neglect all air friction losses.

10.29

15.29

22.94

30.58

Please Explain
Answer #1

M_{1}v_{1}+M_{2}v_{2}= (M_{1}+M_{2})V

V =\frac{M_{1}v_{1}+M_{2}v_{2}}{ (M_{1}+M_{2})}

V =\frac{24.2\times 26.4+44\times 38.5}{ (24.2+44)}= 34.21 mph

V = 34.21 mph = 15.29 m/s

So option 2 is correct.